Suzi and jack have a combined age of 29. Three years ago, triple suzi's age plus double jacks age was 55. How old are suzi and jack?
Answer by Boreal(15235)

S=x
jack=29-x
three years ago S was x-3 and 3(x-3)+2(26-x)=55; 26-x because 3 years ago
3x-9+52-2x=55
x+43=55
x=12 Suzi
Jack is 17
three years ago, S was 9 and jack 14
3*9+2*14=55